OpenStudy (anonymous):

A rectangular counter 3 feet wide and 5 feet long has a circle cut out of it in order to have a sink installed. The circle has a diameter of 18 inches. What is the approximate area of the remaining countertop surface? Use 3.14 for pi.

OpenStudy (carson889):

Area of the counter is 36 inches * 60 inches = 2160 in^2 Area of the sink: d = 18in , r = 9in, A = pi*r^2 = pi *9^2 = 254.47 in^2 Take counter area and subtract sink area = 2160 - 254.47 in^2 = 1905.53 in^2 So the remaining area is approximately 1905.5 in^2

OpenStudy (anonymous):

Choices; I dont see wat one it is 18.9 ft2 15.0 ft2 29.5 ft2 13.2 ft2